# Scoring, Reasoning, and Selecting the Best! Ensembling Large Language Models via a Peer-Review Process

### v1

@@ Abstract
We propose LLM-PeerReview, an unsupervised LLM Ensemble method that selects the most ideal response from multiple LLM-generated candidates for each query, harnessing the collective wisdom of multiple models with diverse strengths. LLM-PeerReview is built on a novel, peer-review-inspired framework that offers a clear and interpretable mechanism, while remaining fully unsupervised for flexible adaptability and generalization. Specifically, it operates in three stages: For scoring, we use the emerging LLM-as-a-Judge technique to evaluate each response by reusing multiple LLMs at hand; For reasoning, we can apply a principled graphical model-based truth inference algorithm or a straightforward averaging strategy to aggregate multiple scores to produce a final score for each response; Finally, the highest-scoring response is selected as the best ensemble output. LLM-PeerReview is conceptually simple and empirically powerful. The two variants of the proposed approach obtain strong results across four datasets, including outperforming the recent advanced model Smoothie-Global by 6.9% and 7.3% points, respectively.
@@ 1 Introduction
Behind this research enthusiasm, we can observe two main points (Chen et al., 2025; Jiang et al., 2023): 1) Persistent performance concerns: Although large language models can be easily deployed for zero-shot or in-context few-shot inference, they still face common performance issues, such as limited accuracy, hallucinations, and misalignment with human goals; 2) The varying strengths and weaknesses of LLMs: These models display significant behavioral differences, primarily driven by variations in their architecture, scale, training data, dictionary, tokenization and methodology. Consequently, their responses to the same prompt often diverge. With the above two points in mind and inspired by the spirit of Ensemble Learning (Dong et al., 2020), it is reasonable to suggest that relying on a single LLM—even one with a high public ranking or other criteria—may not be the optimal strategy for every user query. Instead, it might be more advantageous to simultaneously consider multiple LLM candidates (which are usable out-of-the-box) and leverage their distinct strengths. This concept is the core focus of the burgeoning field of LLM Ensemble (Chen et al., 2025).
@@
As LLM Ensemble gains increasing attention, one well-established class of solutions—ensemble-after-inference (also known as post-hoc ensemble) methods (Jiang et al., 2023; Lv et al., 2024; Tekin et al., 2024; Li et al., 2024c; Guha et al., 2024; Si et al., 2023)—has emerged. These methods include the following two representative approaches (Chen et al., 2025):
@@
  • Selection-then-regeneration approach (Jiang et al., 2023; Lv et al., 2024; Tekin et al., 2024), during inference on a downstream task, first employs a pre-trained “PairRanker” module to select the top-K candidate responses—those deemed most likely to be of high quality—from a pool of LLM-generated responses. This selected subset is then fed into another fine-tuned LLM (e.g., Flan-T5-XL (Chung et al., 2024)) to synthesize a final response. While this line of work has attracted significant attention (Jiang et al., 2023), they rely heavily on carefully curated task-specific training data and the need to fine-tune an additional LLM, limiting their generalization and adaptability.
@@
  • Similarity-based selection approach (Li et al., 2024c; Guha et al., 2024; Si et al., 2023), instead, are mostly fully unsupervised (Guha et al., 2024; Si et al., 2023). These methods follow a simple and intuitive principle: for a given query, select the response with the highest total similarity to all other responses. While such methods pioneered unsupervised post-hoc LLM ensemble, their design remains coarse-grained—they rely on the naive similarity-based selection strategy (Li et al., 2024c; Guha et al., 2024), along with shallow similarity measure of BLEU (Li et al., 2024c) or limited informational utilization (Guha et al., 2024). Thus, the true potential of selection-based post-hoc ensemble remains largely untapped.
@@
When we revisit this research problem, we ask the most fundamental question: In the real world, how would humans select the most ideal text from a set of candidate texts? Perhaps the most immediate and relatable real-world example is: the academic peer-review process.
@@
Motivated by this, we propose a new, fully unsupervised LLM Ensemble method called LLM-PeerReview. Specifically, LLM-PeerReview is structured sequentially around three components: 1) Scoring (analogous to paper reviewing): Given multiple candidate responses to the same query, we adopt the LLM-as-a-Judge approach—leveraging available LLMs as evaluators that assess each response and assign a score (e.g., 5.0 indicating Strong Accept, 4.0 indicating Weak Accept, etc.). To improve scoring accuracy and reduce bias, we further introduce a technique termed “flipped-triple scoring trick” into the Scoring process; 2) Reasoning (analogous to final score estimation made by the senior reviewer): Besides being able to perform direct averaging calculations, we can invoke graphical model-based truth inference techniques from crowdsourcing and weak supervision literature, to perform refined, reliability-aware weighted score aggregation, deriving a final score for each response; 3) Selection (analogous to final decision made by the senior reviewer): This step is analogous to how a senior reviewer or area chair selects the most suitable paper from a small set of submissions. For each query, once final scores have been inferred for all responses, we select the highest-scoring response as the ensemble result.
LLM-PeerReview is built upon the unsupervised, selection-based paradigm, and introduces a novel peer-review-inspired framework for LLM Ensemble, offering a clear and interpretable mechanism. Within the methodology, LLM-PeerReview leverages the emerging LLM-as-a-Judge technique (Li et al., 2024a, b) to evaluate each candidate response, thereby effectively reusing the collective intelligence of multiple existing LLMs at hand. Moreover, the use of a graphical-model-based truth inference algorithm allows us to benefit from the principled graphical model for refined and reliability-aware aggregation of multiple scoring signals. Empirically, extensive experiments conducted with 7B-scale LLMs show that the proposed LLM-PeerReview outperforms recent advanced similarity-based methods (Li et al., 2024c; Guha et al., 2024).
@@ 1.1 Related Work
LLM Ensemble, as outlined in Section 1, can be broadly categorized into three approaches (Chen et al., 2025). The first category, ensemble-before-inference approach (Shnitzer et al., 2023; Srivatsa et al., 2024; Ong et al., 2024), typically necessitates custom-labeled data to pretrain a classifier that routes each query. The mandatory pretraining phase and the dependency on labeled data are the primary inconveniences of these methods. The second category, ensemble-during-inference approach, can be subdivided into token-level (Yu et al., 2024; Huang et al., 2024; Xu et al., 2024), span-level (Liu et al., 2024b; Xu et al., 2025), and process-level (Park et al., 2024) methods, depending on the level of granularity of the information considered in the ensemble. These methods, however, entail substantial computational costs and require the local deployment of every LLM in the ensemble. Lastly, as mentioned before, ensemble-after-inference approach mainly includes selection-then-regeneration-based (Lu et al., 2024) and aggregation-based methods (Li et al., 2024c; Guha et al., 2024).
@@
Figure 1: The proposed LLM-PeerReview contains three steps: (1) Scoring: For a given query, after each LLM independently generates a response (analogous to a submitted academic paper), LLM-PeerReview applies the LLM-as-a-Judge technique (and the proposed flipped-triple scoring trick), treating each model as a reviewer to assign scores to all candidate responses; (2) Reasoning: LLM-PeerReview then uses a truth inference algorithm—analogous to a senior reviewer—to estimate a final score for each response. (Notably, for the variant LLM-PeerReview-Weighted, the inference algorithm is performed using score information across all queries, allowing the model to learn each LLM’s scoring behavior using global information from the dataset, thereby enabling fine-grained, reliability-aware score aggregation); (3) Selecting the best: Finally, for each query, LLM-PeerReview selects the response with the highest final score as the ensemble output—analogous to how a senior reviewer chooses the best paper from a specific submission pool.
@@
Weak Supervision (Zhang et al., 2021; Chen et al., 2023b), also commonly referred to as Learning from Crowds (Chen et al., 2021b, 2022), is a research problem that closely resembles post-inference ensemble methods. The key distinction is that Weak Supervision methods either focus on learning classifiers directly from imperfectly labeled data (Zhang et al., 2021) or on aggregating weak label information (Zheng et al., 2017; Chen et al., 2022), whereas ensemble-after-inference LLM Ensemble methods are exclusively concerned with aggregation and do not involve the learning of classifier models. Additionally, a significant difference is that most weakly supervised learning methods are primarily focused on classification scenarios with closed answer sets, rather than text generation tasks that involve open-ended output spaces.
@@
LLM-as-a-Judge approaches has received a lot of attention recently. As task complexity increases and model outputs become more diverse, traditional evaluation methods—such as matching-based or embedding-based metrics—often fail to capture subtle attributes and provide reliable results (Gu et al., 2024). The recent emergence of large language models has led to the development of the “LLM-as-a-judge” paradigm, where LLMs assess the quality of model outputs. These methods can be broadly classified into three categories: Single-LLM-based, Multi-LLM-based, and Human-AI collaboration-based evaluation approaches. Single-LLM-based methods primarily focus on prompt design (Fu et al., 2023; Kotonya et al., 2023), fine-tuning (Chen et al., 2023a; Wu et al., 2024), or post-processing (Daynauth & Mars, 2024). Notably, Multi-LLM-based methods, which include collaborative (Zhang et al., 2023), competitive (Owens et al., 2024), and aggregation-based (Verga et al., 2024; Chen et al., 2024; Chu et al., 2024) strategies, are closely related to our approach (especially for aggregation methods).
@@ 2 LLM-PeerReview
Without access to any reference responses (i.e., ideal/truth responses), we are given a set of queries $\{\mathbf{x}^{(i)}\}_{i=1}^{I}$ for a generative task. We have access to $J$ large language models $\{\mathcal{M}_{j}\}_{j=1}^{J}$, where each model $\mathcal{M}_{j}$ generates a response $\mathbf{r}^{(i,j)}=\mathcal{M}_{j}(\mathbf{x}^{(i)})$—which is often not ideal—for a given query $\mathbf{x}^{(i)}$. Thus, for each query, we have a set of zero-shot inference responses $\mathbf{R}^{(i)}=[\mathbf{r}^{(i,1)},\ldots,\mathbf{r}^{(i,J)}]$ from heterogeneous LLMs $\{\mathcal{M}_{j}\}_{j=1}^{J}$, while the underlying reference response $\mathbf{y}^{(i)}$ is unobserved to us. Our goal is to ensemble the LLM responses to produce a single, high-quality final response for each query $\mathbf{x}^{(i)}$, using the available data $\mathcal{D}=\{\mathbf{x}^{(i)},\mathbf{R}^{(i)}\}_{i=1}^{I}$.
@@ 2.1 Scoring
Naive point-wise scoring.
As shown in Figure 1 and within our proposed LLM-PeerReview, the scoring phase occurs after the LLMs have first generated responses to the input queries. Using the LLM-as-a-Judge technique (Gu et al., 2024), each LLM judge can assign a point-wise score to each response, representing its overall quality. For example, the score can range from [1, 2, 3, 4, 5], representing the levels of [“Very Poor”, “Poor”, “Acceptable”, “Good”, “Excellent”]. 1111) It is worth noting that each scoring prompt contains both the corresponding query and the response to be evaluated, rather than presenting the response alone; 2) The scoring prompts that we designed and utilized in the experiments are provided in the Appendix.
@@
The above naive point-wise scoring technique can provide scores; however, we propose a technique called flipped triple scoring, which we recommend when applying our approach. Specifically: 1) For multiple responses from different models to the same query $\mathbf{x}^{(i)}$, we first shuffle them; 2) Then, for each LLM judge $\mathcal{M}_{j^{\prime}}$, we score the response triplet $\mathbf{r}^{(i,j-1)},\mathbf{r}^{(i,j)},\mathbf{r}^{(i,j+1)}]$ sequentially (with $J$ times), and for each iteration, we also score the flipped version (i.e., $\mathbf{r}^{(i,j+1)},\mathbf{r}^{(i,j)},\mathbf{r}^{(i,j-1)}]$). As a result, each response receives six scores from the same LLM judge. We can simply average these scores to obtain a final score, which serves as the score $y^{(i,j;j^{\prime})}$ for response $\mathbf{r}^{(i,j)}$ by LLM judge $\mathcal{M}_{j^{\prime}}$. In short, this technique mitigates two common scoring biases in LLM-as-a-Judge (Wang et al., 2023; Zheng et al., 2023; Gu et al., 2024). First, in point-wise scoring, models tend to show a consistent bias toward certain scores (e.g., consistently assigning a score of “1”), as they evaluate a single response without the reference effect of multiple responses. Also, when multiple responses (such as two or three) are presented for evaluation at one time, models frequently exhibit a position bias, tending to favor responses that appear either at the beginning or at the end.
@@ 2.2 Reasoning: a Truth Inference Process
First variant: LLM-PeerReview-Average.
After the scoring phase, as shown in Figure 1, each response $\mathbf{r}^{(i,j)}$ corresponding to a query $\mathbf{x}^{(i)}$ receives multiple scores $\{y^{(i,j;j^{\prime})}\}_{j^{\prime}=1}^{J}$ provided by multiple LLM judges. Then, how can we aggregate these scores meaningfully to compute a final, reliable score $\hat{t}^{(i,j)}$ for each response $\mathbf{r}^{(i,j)}$? This problem is analogous to designing an algorithm that simulates a senior reviewer who consolidates evaluations from multiple reviewers with different scoring preferences and evaluation capabilities. A straightforward and intuitive approach is “averaging” (Zheng et al., 2017; Zhang et al., 2021)—simply taking the mean of all the scores for a given response. As a simple variant of our approach, we refer to it as LLM-PeerReview-Average.
Second variant: LLM-PeerReview-Weighted.
We observe that the above averaging strategy assumes all models to be equally reliable, ignoring the inherent differences in evaluation quality across models. Considering this point, we further propose a weighted variant, referred to as LLM-PeerReview-Weighted. Here, we invoke the well-established Dawid-Skene (DS) model (Dawid & Skene, 1979), a canonical truth-inference graphical model widely used in weak supervision learning and crowdsourcing (Zheng et al., 2017; Zhang et al., 2021), and adapt it to our context. In the following, we introduce the construction of the graphical model in Section 2.2.1, and present the optimization objective and optimization (to obtain the final score value for each response) in Section 2.3.
2.2.1 Graphical Model
Figure 2: Probabilistic graphical representation.
Overall, to infer the underlying “truth” score (unobserved) behind the multiple weak/non-ideal score annotations (observed) for each response, we construct a latent variable graphical model (Bishop & Nasrabadi, 2006; Everett, 2013; Chen et al., 2023b) that includes a latent variable representing the truth score. As depicted the graphical representation in Figure 2, we next introduce the probabilistic generative process we construct from truth scores to weak scores labeled by LLM judges.
First, for each response $\mathbf{r}^{(i,j)}$, we assume that its true score $t^{(i,j)}$ is drawn from a categorical distribution:

### v2

@@ Abstract
We propose LLM-PeerReview, an unsupervised LLM Ensemble method that selects the most ideal response from multiple LLM-generated candidates for each query, harnessing the collective wisdom of multiple models with diverse strengths. LLM-PeerReview is built on a novel, peer-review-inspired framework that offers a transparent and interpretable mechanism, while remaining fully unsupervised for flexible adaptability and generalization. Specifically, it operates in three stages: For scoring, we use the emerging LLM-as-a-Judge technique to evaluate each response by reusing multiple LLMs at hand; For reasoning, we can apply a straightforward averaging strategy or a principled graphical model-based truth inference algorithm to aggregate multiple scores to produce a final score for each response; Finally, the highest-scoring response is selected as the best ensemble output. LLM-PeerReview is conceptually simple and empirically powerful. Our results across four datasets show that the two variants of the proposed approach outperform the advanced model Smoothie-Global by 6.9% and 7.3% points, cross diverse task types including factual recall QA, math reasoning, and instruction following.
@@ 1 Introduction
Behind this research enthusiasm, we can observe two main points (Jiang et al., 2023): 1) Persistent performance concerns: Although large language models can be easily deployed for zero-shot or in-context few-shot inference, they still face common performance issues, such as limited accuracy, hallucinations, and misalignment with human goals; 2) The varying strengths and weaknesses of LLMs: These models display significant behavioral differences, primarily driven by variations in their architecture, scale, training data, dictionary, tokenization and methodology. Consequently, their responses to the same prompt often diverge. With the above two points in mind and inspired by the spirit of Ensemble Learning (Dong et al., 2020), it is reasonable to suggest that it is advantageous to simultaneously consider multiple LLMs (which are usable out-of-the-box) and leverage their distinct strengths. This concept is the core focus of the burgeoning field of LLM Ensemble (Chen et al., 2025b).
@@
As LLM Ensemble gains increasing attention, one well-established class of solutions—ensemble-after-inference (also known as post-hoc ensemble) methods—has emerged. These methods include the following two categories:
@@
  • Selection-then-regeneration approach (Jiang et al., 2023; Lv et al., 2024; Tekin et al., 2024), first employs a pre-trained “PairRanker” module to select the top-K candidate responses—those deemed most likely to be of high quality—from a pool of LLM-generated responses. This selected subset is then fed into another fine-tuned LLM (e.g., Flan-T5-XL (Chung et al., 2024)) to synthesize a final response. While these methods have attracted significant attention (Jiang et al., 2023), they rely on carefully curated task-specific training data and the need to fine-tune an additional LLM, limiting their generalization and adaptability.
@@
  • Similarity-based selection approach (Li et al., 2024; Guha et al., 2024; Si et al., 2023), instead, is mostly fully unsupervised (Guha et al., 2024; Si et al., 2023). These methods follow a simple principle: for a given query, select the response with the highest total similarity to all other responses. While such methods pioneered unsupervised post-hoc LLM Ensemble, their design remains coarse-grained—they rely on the naive similarity-based selection strategy (Li et al., 2024; Guha et al., 2024), along with shallow similarity measure of BLEU (Li et al., 2024) and limited informational utilization (Guha et al., 2024). Thus, the true potential of the ensemble remains largely untapped.
@@
When we revisit this research problem, we ask the most fundamental question: In the real world, how would humans select the most ideal text from a set of candidate texts? Perhaps the most immediate and relatable real-world example is: the academic peer-review process. Motivated by this, we propose a new, fully unsupervised LLM Ensemble approach called LLM-PeerReview.
@@
Specifically, LLM-PeerReview comprises three sequential modules: (1) Scoring (analogous to paper reviewing): Given multiple candidate responses to the same query, we adopt the LLM-as-a-Judge approach—leveraging available LLMs as evaluators that assess each response and assign a score (e.g., 5.0 indicating Strong Accept, 4.0 indicating Weak Accept, etc.). To enhance scoring precision and mitigate inherent biases, we further propose the flipped-triple scoring trick—a cornerstone mechanism that fundamentally determines the efficacy of the entire framework; (2) Reasoning (analogous to final score estimation made by the senior reviewer): Besides being able to perform direct averaging calculations, we can invoke graphical model-based truth inference techniques from weak supervision literature, to perform reliability-aware weighted score aggregation, deriving a final score for each response; (3) Selection (analogous to final decision made by the senior reviewer): This step is analogous to how a senior reviewer or area chair selects the most suitable paper from a specific submission pool. For each query, we identify the highest-scoring response as the final ensemble result once all scores have been inferred.
LLM-PeerReview is built upon a transparent and interpretable framework that mimics the peer-review process, where its fully unsupervised nature provides flexibility and generalizability across various tasks and datasets. Within the methodology, we employ the LLM-as-a-Judge approach to leverage collective intelligence of LLMs for refined evaluation, moving beyond coarse metrics like BLEU, and propose a novel pivotal flipped-triple scoring trick to reduce scoring bias. Further, we provide both straightforward averaging and graphical model-based weighted methods to aggregate multiple scoring signals. Extensive experiments show that the proposed LLM-PeerReview significantly outperforms recent advanced similarity-based methods (Li et al., 2024; Guha et al., 2024) and all individual LLMs.
@@ 1.1 Related Work
LLM Ensemble, can be broadly categorized into three approaches. The first category, ensemble-before-inference approach (Shnitzer et al., 2023; Srivatsa et al., 2024; Ong et al., 2024), typically necessitates custom-labeled data to pretrain a classifier that routes each query. The mandatory pretraining phase and the dependency on labeled data are the primary inconveniences of these methods. The second category, ensemble-during-inference approach, can be subdivided into token-level (Yu et al., 2024; Huang et al., 2024; Xu et al., 2024; Zou et al., 2025; Chen et al., 2025a), span-level (Liu et al., 2024b; Xu et al., 2025), and process-level (Park et al., 2024) methods, depending on the level of granularity of the information considered in the ensemble. Among them, some methods (Zou et al., 2025; Chen et al., 2025a) require post-training (Yang et al., 2026). Lastly, as mentioned before, ensemble-after-inference approach mainly includes selection-then-regeneration-based (Lu et al., 2024) and aggregation-based methods (Li et al., 2024; Guha et al., 2024). Also, there are several approaches closely related to LLM Ensemble, such as using multiple prompts for ensemble inference (He et al., 2025).
@@
Figure 1: The proposed LLM-PeerReview contains three steps: (1) Scoring: For a given query, after each LLM independently generates a response (analogous to a submitted academic paper), LLM-PeerReview applies the LLM-as-a-Judge technique (and the proposed flipped-triple scoring trick), treating each model as a reviewer to assign scores to all candidate responses; (2) Reasoning: LLM-PeerReview then uses a truth inference algorithm—analogous to a senior reviewer—to estimate a final score for each response. (Notably, for the variant LLM-PeerReview-W, the inference algorithm is performed using score information across all queries, allowing the model to learn each LLM’s scoring behavior using global information from the dataset); (3) Selecting the best: Finally, for each query, LLM-PeerReview selects the response with the highest final score as the ensemble output—analogous to how a senior reviewer chooses the most ideal paper from a specific submission pool.
@@
Weak Supervision (Zhang et al., 2021; Chen et al., 2023b), also commonly referred to as Learning from Crowds (Chen et al., 2021b, 2022), is a research problem that closely resembles post-inference ensemble methods. The key distinction is that Weak Supervision methods either focus on learning classifiers directly from imperfectly labeled data (Zhang et al., 2021) or on aggregating weak label information (Zheng et al., 2017; Chen et al., 2022), whereas ensemble-after-inference LLM Ensemble methods are exclusively concerned with aggregation and do not involve the learning of classifier models. Additionally, a significant difference is that most weakly supervised learning methods are primarily focused on classification with closed answer sets, rather than open-ended generation tasks.
@@
LLM-as-a-Judge approaches has received a lot of attention recently. As task complexity increases and model outputs become more diverse, traditional evaluation methods—such as matching-based or embedding-based metrics—often fail to capture subtle attributes. The recent emergence of LLMs has led to the “LLM-as-a-judge” paradigm, where LLMs assess the quality of model outputs. These methods can be broadly classified into three categories: Single-LLM-based, Multi-LLM-based, and Human-AI collaboration-based evaluation approaches. Single-LLM-based methods primarily focus on prompt design (Fu et al., 2023), fine-tuning (Chen et al., 2023a), or post-processing (Daynauth and Mars, 2024). Notably, Multi-LLM-based methods, which include collaborative (Zhang et al., 2023), competitive (Owens et al., 2024), and aggregation-based (Verga et al., 2024; Chen et al., 2024) strategies, are closely related to our approach.
@@ 2 LLM-PeerReview
Without access to any reference responses (i.e., ideal/truth responses), we are given a set of queries $\{\mathbf{x}^{(i)}\}_{i=1}^{I}$. We have access to $J$ large language models $\{\mathcal{M}_{j}\}_{j=1}^{J}$, where each model $\mathcal{M}_{j}$ generates a response $\mathbf{r}^{(i,j)}=\mathcal{M}_{j}(\mathbf{x}^{(i)})$—which is often not ideal—for a given query $\mathbf{x}^{(i)}$. Thus, for each query, we have a set of zero-shot inference responses $\mathbf{R}^{(i)}=[\mathbf{r}^{(i,1)},\ldots,\mathbf{r}^{(i,J)}]$ from heterogeneous LLMs $\{\mathcal{M}_{j}\}_{j=1}^{J}$, while the underlying reference response $\mathbf{y}^{(i)}$ is unobserved to us. Our goal is to ensemble the LLM responses to produce a single, high-quality final response for each query $\mathbf{x}^{(i)}$, using the available data $\mathcal{D}=\{\mathbf{x}^{(i)},\mathbf{R}^{(i)}\}_{i=1}^{I}$.
@@ 2.1 Scoring
As shown in Figure 1 and within our proposed LLM-PeerReview, the scoring phase occurs after the LLMs have first generated responses to the input queries. Using the LLM-as-a-Judge technique (Gu et al., 2024), each LLM judge can assign a point-wise score to each response, representing its overall quality.   [^1] For example, the score can range from [1, 2, 3, 4, 5], representing the levels of [“Very Poor”, “Poor”, “Acceptable”, “Good”, “Excellent”].
@@
The above naive point-wise scoring technique can provide scores; however, we propose a new technique called flipped triple scoring, which we recommend when applying our approach.  [^2] It is worth noting that each scoring prompt contains both the corresponding query and the response to be evaluated, rather than presenting the response alone. The scoring prompts that we designed and utilized in the experiments are provided in the Appendix. Specifically: 1) For multiple responses from different models to the same query $\mathbf{x}^{(i)}$, we first shuffle them; 2) Then, for each LLM judge $\mathcal{M}_{j^{\prime}}$, we score the response triplet $\mathbf{r}^{(i,j-1)},\mathbf{r}^{(i,j)},\mathbf{r}^{(i,j+1)}]$ sequentially (with $J$ times), and for each iteration, we also score the flipped version (i.e., $\mathbf{r}^{(i,j+1)},\mathbf{r}^{(i,j)},\mathbf{r}^{(i,j-1)}]$). As a result, each response receives six scores from the same LLM judge. We can simply average these scores to obtain a final score, which serves as the score $y^{(i,j;j^{\prime})}$ for response $\mathbf{r}^{(i,j)}$ by LLM judge $\mathcal{M}_{j^{\prime}}$. In short, this technique mitigates two common scoring biases in LLM-as-a-Judge (Wang et al., 2023; Zheng et al., 2023; Gu et al., 2024). First, in point-wise scoring, models tend to show a consistent bias toward certain scores (e.g., consistently assigning a score of “1”), as they evaluate a single response without the reference effect of multiple responses. Also, when multiple responses (such as two or three) are presented for evaluation at one time, models frequently exhibit a position bias, tending to favor responses that appear either at the beginning or at the end.
@@ 2.2 Reasoning: a Truth Inference Process
For the Reasoning process within LLM-PeerReview, as noted in Section  1 and the caption of Figure 1, we can adopt different specific sub-components: 1) A default, straightforward, yet effective averaging method, as described in Section  2.2.1; 2) A weighted method, LLM-PeerReview-W, as detailed in Section 2.2.2.
2.2.1 Reasoning for Default LLM-PeerReview
After the scoring phase, as shown in Figure 1, each response $\mathbf{r}^{(i,j)}$ corresponding to a query $\mathbf{x}^{(i)}$ receives multiple scores $\{y^{(i,j;j^{\prime})}\}_{j^{\prime}=1}^{J}$ provided by multiple LLM judges. Then, how can we aggregate these scores meaningfully to compute a final, reliable score $Score(\mathbf{r}^{(i,j)})$ for each response $\mathbf{r}^{(i,j)}$, using scores $\{y^{(i,j;j^{\prime})}\}_{j^{\prime}=1}^{J}$? This problem is analogous to designing an algorithm that simulates a senior reviewer who consolidates evaluations from multiple reviewers with different scoring preferences and evaluation capabilities. A straightforward and intuitive approach is averaging (Zheng et al., 2017; Zhang et al., 2021)—simply taking the mean of all the scores for a given response:
$Score(\mathbf{r}^{(i,j)})=\frac{1}{J}\sum_{j^{\prime}=1}^{J}y^{(i,j;j^{\prime})}.$ (1)
2.2.2 Reasoning for LLM-PeerReview-W
We observe that the above averaging strategy assumes all models to be equally reliable. We further propose a weighted variant, referred to as LLM-PeerReview-W. We invoke the well-established Dawid-Skene (DS) model (Dawid and Skene, 1979), a canonical truth-inference graphical model widely used in weak supervision learning  (Zhang et al., 2021). We implement several necessary improvements to the origin model, particularly regarding the processing of regression scores. In the following, we introduce the construction of the graphical model, and present the optimization objective and optimization (to obtain the final score $Score(\mathbf{r}^{(i,j)})$ for each response $\mathbf{r}^{(i,j)}$ for each response $\mathbf{r}^{(i,j)}$). In brief and intuitively, we perform two iterative steps: (1) inferring the true score category for each response, and (2) learning the parameters of the graphical model.
Graphical model construction.
Overall, to infer the underlying “truth” score (unobserved) behind the multiple weak/non-ideal score annotations (observed) for each response, we construct a latent variable graphical model (Bishop and Nasrabadi, 2006; Everett, 2013) that includes a latent variable representing the truth score. As depicted the graphical representation in Figure 2, for each response $\mathbf{r}^{(i,j)}$, we assume that its true score $t^{(i,j)}$ is drawn from a categorical distribution: